OpenStudy (anonymous):

i need to fin the smallest possible surface area of any 3D shape but it has to have a volume of 1200cm3

OpenStudy (anonymous):

Spheres have the smallest surface area to volume ratio. Volume of a sphere=(4/3)pi*r^3 plug in 1200 1200=(4/3)pi*r^3 Find the radius by solving for r Surface Area of a Sphere=4*pi*r^2 Once you've found r, plug it into this equation to find the corresponding surface area.